the second side of a triangle is twice the first side, and the third is 6 less than the second side. the perimeter is 94 ft. find the sides
I tried x+2x+2x-6 =5x-6=94
95/5x=18.8 x 2 =37.6 -6 =88
sides = x=18 2x=37.6 and 2x-6=31.6 perimeter =88 where did i go wrong?

Answer by jonvaliente(64)

let 2x=length of second side
x=length of first side
2x-6=length of third side
add them all up to get the perimeter
2x+x+2x-6=94
5x-6=94
5x=100
x=20
first side = 20
second side =40
third side = 34